# Human-in-the-Loop Causal Discovery under Latent Confounding using Ancestral GFlowNets

### v2

@@ 5.1 Distributional Assessment of AGFN
Results. Figure 3 shows that the AGFN adequately approximates the theoretical distribution induced by the reward in Equation 9. Furthermore, AGFNs induce distributions over BIC and SHD values that closely resemble those induced by $p(\mathcal{G})\propto R(\mathcal{G})$. We also note an important improvement over the prior art on probabilistic CD (N-ADMG): we found that over $60\%$ of its samples were non-ancestral and that this method was of little use for making inferences over AGs. Meanwhile, AGFN does not sample non-ancestral graphs.
5.2 Comparison with SOTA CD algorithms
(a) chain4
(b) IV
(c) collfork
Figure 4: Ancestral graphs representing the data generating models for the three considered datasets in Table 1.
Data. We generate $10$ datasets with $500$ independent samples from the randomly parametrized linear Gaussian SCMs corresponding to the canonical causal diagram Richardson and Spirtes (2002) in each AG depicted in Figure 4. Unshielded colliders and discriminating paths are fundamental patterns in the detection of invariances by CD algorithms under latent confounding Spirtes and Richardson (1997); Zhang (2008b). Thus, we consider the following 4-node causal diagrams with increasingly difficult configurations: (i) chain4, a chain without latent confounders; (ii) collfork, a graph with triplets involving colliders and non-colliders under latent confounding, and (iii) IV, a structure with a discriminating path for $Z$: $W\rightarrow X\leftarrow Z\rightarrow Y$.
Baselines. We compare AGFN with four notable CD methods: FCI (Spirtes et al., 2001; Zhang, 2008b), GFCI (Ogarrio et al., 2016), DCD (Bhattacharya et al., 2021), and N-ADMG Ashman et al. (2023). The baselines span four broad classes of CD methods. FCI is a seminal constraint-based CD algorithm that learns a PAG consistent with conditional independencies entailed by statistical tests. GFCI is a hybrid CD algorithm that learns a PAG by first obtaining an approximate structure using FGS (Ramsey, 2015) (a BIC-score-based search algorithm for causally sufficient scenarios) and then by applying FCI to identify possible confounding and remove some edges added by FGS. DCD casts CD as continuous optimization with differentiable algebraic constraints defining the space of AGs and uses gradient-based algorithms to solve it. N-ADMG computes a variational approximation of the joint posterior distribution over the space of bow-free causal diagrams (Nowzohour et al., 2017) associated with non-linear SCMs with additive noise. While N-ADMG focuses on a more restricted setting compared to AGs, it offers some uncertainty quantification in the variational pohttps://www.overleaf.com/project/650454a084b798332af29ebesterior, making it more closely comparable to our approach. We rigorously follow the experimental guidelines in the original works.
Experimental setup We train AGFN on each dataset and use it to sample 100k graphs. We also apply FCI, GFCI, and DCD to $100$ bootstrapped resamplings of each dataset to emulate confidence distributions induced by these algorithms. To compare the algorithms’ outputs, we compute the sample mean and standard deviation of the BIC and SHD at the PAG level. Specifically, we compute the SHD between the ground-truth PAG and each estimated PAG obtained by each method. If the output is a PAG member (as for DCD, N-ADMG, and AGFN) we use FCI to transform the output using these graphs as oracles for conditional independencies. Furthermore, we directly compute the BIC for the outputs, as all PAG members are asymptotically score-equivalent.
Results. Table 1 compares AGFN against baseline CD algorithms. Notably, our method consistently outperforms the only probabilistic baseline in the literature (N-ADMG) in terms of both SHD and BIC. As expected, however, the average BIC and SHD induced by AGFN are larger than those induced by the bootstrapped versions of the non-probabilistic algorithms, and the variances are greater; this is due to the inherent sampling diversity of our method and the resulting generation of possibly implausible samples. Indeed, Table 2 shows that the three most rewarding samples from AGFN are as good as (and sometimes better than) the other CD algorithms. Results for N-ADMG comprise the three most frequent samples from the variational distribution.
chain4 IV collfork
FCI $2.07{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.00}}$ 2.00 $3.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.90}}$ 2.90 $5.43{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.87}}$ 1.87
GFCI $\mathbf{1.50}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.63}}$ 1.63 $3.63{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 3.16}}$ 3.16 $5.53{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.11}}$ 2.11
DCD $2.27{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.46}}$ 1.46 $4.80{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.17}}$ 2.17 $5.60{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\pm\scriptstyle 2.13}}$ 2.13
N-ADMG (top 3) $4.38{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.81}}$ 0.81 $6.08{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.77}}$ 1.77 $6.87{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 0.93}}$ 0.93
AGFN (top 3) $2.00{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 1.55}}$ 1.55 $\mathbf{3.50}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 3.29}}$ 3.29 $\mathbf{4.90}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\scriptstyle\pm 2.70}}$ 2.70
Table 2: SHD for point estimates. The mean SHD of the top-3 AGFN draws is comparable to or better than baselines.
Figure 5: CD with simulated human feedback. The top/bottom row shows the mean SHD/BIC of AGFN samples as a function of human interactions. Probing the expert about the edge that minimizes the mean cross-entropy leads to a faster decrease in BIC compared to a random strategy. The SHD decreases similarly in both cases. Results reflect the outcomes of 30 simulations.
5.3 Simulating humans in the loop
@@
Data. We follow the procedure from Section 5.1 to generate graphs with $4$, $6$, $8$ and $10$ nodes. We draw $500$ samples from a compatible linear Gaussian SCM and use them to train an AGFN. Then, we follow our active elicitation strategy from Section 4 to probe simulated humans, adhering to the generative model described in the same section, with $\pi=0.9$.
@@
Setup. Since we are the first to propose an optimal design for expert knowledge elicitation, there are no baselines to compare AGFN against. That being said, we aim to determine whether the inclusion of expert feedback enhances the concentration of the learned distribution around the true AG, and evaluate the effectiveness of our elicitation strategy. To do so, we measure SHD to the true AG and BIC as a function of the number of expert interactions.
@@
Results. Figure 5 shows that incorporating expert feedback substantially decreases the expected SHD and BIC under our belief over AGs. On the one hand, the remarkable decrease in expected SHD shows that our belief becomes increasingly focused on the true AG as we iteratively request the expert’s feedback, regardless of the querying strategy. On the other hand, the second row shows that our querying strategy results in a substantial decrease in the BIC, demonstrating a faster reduction than random queries. This validates the notion that some edges are more informative than others, and we should prioritize them when probing the expert.
6 Related Work
CD under latent confounding. Following the seminal works by Spirtes et al. (2001) and Zhang (2008b) introducing the complete FCI, a variety of works have emerged. Among them are algorithms designed for sparse scenarios, including RFCI (Colombo et al., 2012) and others (Silva, 2013; Claassen et al., 2013). Notably, Silva (2013)’s framework uses a Bayesian approach to CD of Gaussian causal diagrams based on sparse covariance matrices. Nonetheless, it requires sampling one edge at a time and relies on numerical heuristics that might effectively alter the posterior we are sampling from. Colombo et al. (2012) introduced the conservative FCI to handle conflicts arising from statistical errors in scenarios with limited data, even though it yields less informative results. Subsequent efforts to improve reliability led to the emergence of constraint-based CD algorithms based in Boolean satisfiability (Hyttinen et al., 2014; Magliacane et al., 2016), although they are known to scale poorly on $|\mathbf{V}|$ (Lu et al., 2021). In another paradigm, score-based search algorithms rank MAGs according to goodness-of-fit measures, commonly using BIC for linear Gaussian SCMs (Triantafillou and Tsamardinos, 2016; Zhalama et al., 2017a; Rantanen et al., 2021). There are also hybrid approaches that combine constraint-based strategies to reduce the search space, such as GFCI (Ogarrio et al., 2016), M3HC (Tsirlis et al., 2018), BCCD (Claassen and Heskes, 2012), and GSPo (Bernstein et al., 2020). Continuous optimization approaches have recently emerged as a novel approach to score-based CD, such as DCD (Bhattacharya et al., 2021) and N-ADMG Ashman et al. (2023).
CD with expert knowledge. Previous works on CD have explored various forms of background knowledge. This includes knowledge on edge existence/non-existence Meek (1995b), ancestral constraints Chen et al. (2016), variable grouping Parviainen and Kaski (2017), partial order Andrews (2020) and typing of variables (Brouillard et al., 2022). Incorporating expert knowledge is pivotal to reducing the search space and the size of the learned equivalence class. However, due to significant challenges, up to date, there are only a few works trying to integrate human knowledge into CD within the context of latent confounding (Andrews, 2020; Wang et al., 2022). These works operate under the assumption of perfect expert feedback. In contrast, our contribution is novel in that it confronts the challenges of real-world situations where expert input might be inaccurate.
@@ 7 Discussion
We presented AGFN, the first probabilistic CD method that accounts for latent confounding and incorporates potentially noisy human feedback in the loop. AGFN samples AGs according to a score function, quantifying the uncertainty in the learning process. Furthermore, it can leverage human feedback in an optimal design strategy, efficiently reducing our uncertainty on the true data-generating model.
This work is focused on linear Gaussian models, using BIC as our score. However, the implementation of AGFNs is not restricted by this choice. In principle, we could replace the BIC with alternative score functions that are more appropriate for different types of variables, e.g. for discrete data (Drton and Richardson, 2008). It is also important to highlight that our framework does not require retraining the AGFN after we see human feedback. Moreover, AGFN is a GPU-powered algorithm, and while we used only one GPU in our experiments, it is possible to greatly accelerate AGFN by using cluster architectures with multiple GPUs.
@@
By offering uncertainty-quantified CD together with a recipe for including humans in the loop, we expect AGFNs will significantly enhance the accuracy and reliability of CD, especially in real-world domains. Moreover, AGFNs bring a novel perspective to developing more comprehensive tools for downstream causal tasks Bareinboim and Pearl (2016), as the resulting distribution encodes knowledge from data and human feedback while accounting for epistemic uncertainty. For example, methods for causal reasoning that currently rely on a single AG Zhang (2008a); Jaber et al. (2022) could exploit this distribution to incorporate a richer understanding of uncertainty and knowledge, thereby enhancing their robustness and reliability.
Acknowledgments
@@
Diego Mesquita acknowledges the support by the Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro FAPERJ (SEI-260003/000709/2023), the São Paulo Research Foundation FAPESP (2023/00815-6), the Conselho Nacional de Desenvolvimento Científico e Tecnológico CNPq (404336/2023-0), and the Silicon Valley Community Foundation through the University Blockchain Research Initiative (Grant #2022-199610). António Góis acknowledges the support by Samsung Electronics Co., Ldt. Adèle Ribeiro and Dominik Heider were supported by the LOEWE program of the State of Hesse (Germany) in the Diffusible Signals research cluster and by the German Federal Ministry of Education and Research (BMBF) [031L0267A] (Deep Insight). Samuel Kaski was supported by the Academy of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI), EU Horizon 2020 (European Network of AI Excellence Centres ELISE, grant agreement 951847), UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1). We also acknowledge the computational resources provided by the Aalto Science-IT Project from Computer Science IT.
Appendix A Additional Related Works
Generative Flow Networks (GFlowNets; Bengio et al., 2021a, b) are generative models that sample discrete composite objects from an unnormalized reward function. They have been successfully used to sample various structures such as protein sequences (Jain et al., 2022) and schedules (Zhang et al., 2023). They have also been used to train energy-based models (Zhang et al., 2022). In the field of structure learning, they have been applied to Bayesian networks — more specifically to sample a posterior over DAGs in linear Gaussian networks, although without accounting for unobserved confounding (Deleu et al., 2022). Recently, Deleu et al. (2023) proposed an extension to jointly infer the structure and parameters, also grounded in the assumption of causal sufficiency. It is worth highlighting that training GFlowNets in these scenarios presents optimization challenges, resulting in the utilization of a variety of loss functions (Shen et al., 2023). Moreover, Lahlou et al. (2023) proposed an extension of GFlowNets to continuous domains.
Appendix B Cross-entropy acquisition
@@
The expected mutual information and the information gain are the most widely used information-theoretic measures to actively interact with a human and choose the most informative data points to be labeled Ryan et al. (2015). However, we instead use the negative expected cross-entropy between the current and updated beliefs as the acquisition function of our experimental design (see eq. 17). As we show next, the approximation of both the mutual information and the information gain is intrinsically dependent upon the estimation of the log-partition of the updated beliefs over the space of ancestral graphs. Doing so is computationally intensive, and we would either need to use a Monte Carlo estimator of the integrals or use some posterior approximation — in both cases, leading to asymptotically biased estimates of the acquisition. In contrast, we can easily leverage AGFN samples to compute asymptotically unbiased estimates of our acquisition function. The next paragraphs provide further details.
Mutual information.
@@
as the partition function of our updated beliefs. Note also that Equation 21 entails computing the entropy of $q(\mathcal{G},f_{r}|\mathbf{f}_{K})$. Thus, the selection criterion in eq. 20 requires an accurate estimate of $\log c_{K+1}(f_{r})$ — which is well-known for being a difficult problem Ma et al. (2013) — and the Monte Carlo estimator for the log-partition function is asymptotically biased.
Information gain.
The expected information gain of an elicitation is defined as the expected KL divergence between our updated and current beliefs over ancestral graphs. This approach is widely employed in Bayesian experimental design Ryan et al. (2015). In our framework, the information gain resulting from a feedback $f_{r}$ is
@@
with $f_{r_{K+1}}=f_{r}$, $c_{K}$ as the partition function of $q(\cdot|\mathbf{f}_{K})$ — that does not depend upon $f_{r}$ —, and $c_{K+1}(f_{r})$ defined in eq. 22, the estimation of the information gain is inherently dependent upon the estimation of the log-partition function.
Cross-entropy.

### v3

@@ 5.2 Comparison with SOTA CD algorithms
[…]
5.3 Simulating experts in the loop
@@
Data. Following the procedure from Section 5.1, we generate graphs with $4$, 66, 88 and $10$ nodes and draw $500$ samples from a compatible linear Gaussian SCM to train AGFN. We then apply our active elicitation strategy (Alg. 1) to probe simulated experts, adhering to the generative model in Section 4, with $\pi=0.9$.
@@
Setup. Since benchmarks for expert knowledge elicitation are lacking, we assess whether incorporating expert feedback enhances the concentration of the learned distribution around the true AG and evaluate our elicitation strategy by tracking SHD and BIC as a function of the number of expert interactions.
@@
Results. Figures 4 and 5 show that the expected SHD under our belief over AGs substantially decreases with expert feedback. The consistent decrease in SHD indicates that our belief increasingly concentrates on the true AG as feedback is iteratively incorporated, regardless of the querying strategy. Moreover, our querying strategy accelerates BIC reduction compared to random queries (see Section C.8), confirming that some edges are more informative and should be prioritized when querying the expert.
Expert-aided AGFN versus baselines. Figure 4 indicates that updating AGFN’s distribution by incorporating the expert’s feedback leads to a significant enhancement of the model’s accuracy in finding the underlying causal diagram (ACI is omitted due to its significant lower performance). Remarkably, even a single feedback is sufficient for AGFN to outperform all considered baselines. Importantly, integrating expert knowledge into prior CD methods is generally nontrivial: when feasible, it typically occurs before inference (Andrews,, 2020) or assumes perfect rather than uncertain knowledge (Wang et al.,, 2022).
5.4 GPT-4o as the expert-in-the-loop
Figure 5: GPT-aided AGFN outperforms CD baselines in the considered datasets. Each plot shows the expected SHD under a varying number of feedbacks over 33 EITL simulations.
Data. We use three variants of the well-known Sachs dataset (Sachs et al.,, 2005), namely Sachs-5 with 5 variables, Sachs-7 with 7 variables, and the original Sachs with 11 variables, to analyze CD under latent confounding. In each variant, the variables with the most false observed independencies were made latent, ensuring that the resulting graphs remain sufficiently connected. Further details are in Section C.3.
Setup. We again apply our active elicitation strategy as described in Section 4 with few adjustments. A complete description of the LLM elicitation framework can be found in Section D.1. We prompt the LLM using simple strategies similar to  Kiciman et al., (2024); Bazaluk et al., (2025). For these experiments, each GPT-4o (OpenAI et al.,, 2024) answer is assigned a distinct reliability $\pi$. There are many works which discuss about deriving uncertainty over the response of an LLM (Xiong et al.,, 2024; Tanneru et al.,, 2024; Ma et al.,, 2025). In this work, we use a Monte Carlo estimate with 10 sampled GPT-4o generations for each query (Bazaluk et al.,, 2025); see Appendix D.
Results. Figure 5 shows that incorporating GPT-4o as an expert leads to a consistent reduction in expected SHD as more feedback is provided. Although the LLM is not a perfect expert, occasionally providing incorrect responses, the results exhibit robust performance. This highlights that explicitly accounting for varying levels of uncertainty across different LLM responses can substantially improve the accuracy and reliability of our model, even when expert feedback is imperfect.
@@ 6 DISCUSSION
We presented AGFN, the first probabilistic CD method that accounts for latent confounding and incorporates potentially noisy expert feedback in the loop. AGFN samples AGs according to a score function, quantifying the uncertainty in the learning process. Furthermore, it can leverage expert feedback in an optimal design strategy, efficiently reducing our uncertainty on the true data-generating model.
Notably, AGFNs are not limited to a specific choice of score. In principle, the BIC could be replaced with alternative scoring functions better suited for different types of variables, such as discrete data (Drton and Richardson,, 2008). Furthermore, our framework does not require retraining the AGFN after incorporating expert feedback, allowing efficient updates.
By combining uncertainty-quantified CD with a systematic approach for incorporating experts in the loop, AGFNs are expected to substantially improve the accuracy and reliability of CD, particularly in real-world domains. Our results further demonstrate that, even when using imperfect experts such as an LLM, the framework successfully converges to the true AG.
@@
Beyond this, AGFNs offer a novel perspective for developing more comprehensive tools for downstream causal tasks (Bareinboim and Pearl,, 2016), as the resulting distribution integrates both data-driven and expert-informed knowledge while explicitly accounting for epistemic uncertainty. For instance, causal reasoning methods that traditionally rely on a single AG (Zhang, 2008a, ; Jaber et al.,, 2022) could leverage this distribution to incorporate richer uncertainty estimates and expert knowledge, thereby enhancing robustness and reliability.
Acknowledgements
@@
Tiago da Silva, Eliezer Silva and Diego Mesquita acknowledge the support of Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro FAPERJ (SEI-260003/000709/2023), São Paulo Research Foundation FAPESP (2023/00815-6), Conselho Nacional de Desenvolvimento Científico e Tecnológico CNPQ (404336/2023-0, 445170/2024-7, 305692/2025-9). Bruna Bazaluk acknowledges the support of Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Código de Financiamento 001. Adèle Ribeiro and Dominik Heider were supported by the LOEWE program of the State of Hesse (Germany) in the Diffusible Signals research cluster and by the German Federal Ministry of Research, Technology and Space (BMBTR) [031L0267A] (Deep Insight). António Góis acknowledges the Canada CIFAR AI Chair program (Mila), by an unrestricted gift from Google, and by Samsung Electronics Co., Ldt. Samuel Kaski was supported by the Academy of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI), EU Horizon 2020 (European Network of AI Excellence Centres ELISE, grant agreement 951847), UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1). We also acknowledge the computational resources provided i) by the Aalto Science-IT Project from Computer Science IT, and ii) FGV TIC.
Appendix A Cross-entropy acquisition
@@
The expected mutual information and the information gain are the most widely used information-theoretic measures to actively interact with a human and choose the most informative data points to be labeled (Ryan et al.,, 2015). However, we instead use the negative expected cross-entropy between the current and updated beliefs as the acquisition function of our experimental design (see eq. 13). As we show next, the approximation of both the mutual information and the information gain is intrinsically dependent upon the estimation of the log-partition of the updated beliefs over the space of ancestral graphs. Doing so is computationally intensive, and we would either need to use a Monte Carlo estimator of the integrals or use some posterior approximation – in both cases, leading to asymptotically biased estimates of the acquisition. In contrast, we can easily leverage AGFN samples to compute asymptotically unbiased estimates of our acquisition function. The next paragraphs provide further details.
A.1 Mutual information
@@
as the partition function of our updated beliefs. Note also that Equation 17 entails computing the entropy of $q(\mathcal{G},f_{r}|\mathbf{f}_{K})$. Thus, the selection criterion in eq. 16 requires an accurate estimate of $\log c_{K+1}(f_{r})$, which is well-known for being a difficult problem (Ma et al.,, 2013), and the Monte Carlo estimator for the log-partition function is asymptotically biased.
A.2 Information gain
The expected information gain of an elicitation is defined as the expected KL divergence between our updated and current beliefs over ancestral graphs. This approach is widely employed in Bayesian experimental design (Ryan et al.,, 2015). In our framework, the information gain resulting from a feedback $f_{r}$ is
@@ A.2 Information gain
with $f_{r_{K+1}}=f_{r}$, $c_{K}$ as the partition function of $q(\cdot|\mathbf{f}_{K})$ that does not depend upon $f_{r}$, and $c_{K+1}(f_{r})$ defined in eq. 18, the estimation of the information gain is inherently dependent upon the estimation of the log-partition function.
A.3 Cross-entropy